If you had a 1 cubic inch closed container filled with Co2 heated up to 700C and also a 1 cubic inch closed container filled with helium heated up to 700C would the pressure be the same in both chambers or differ, and why?
Answer #1

The ideal gas equation is given by:

PV = nRT

Where

P- pressure

V- volume

n- Number of moles of the gas atom

R- universal gas constant = 8.314 J/mol.K.

T- temperature in Kelvin

For CO2

V = 1 cubic inch = 1.639 × 10-5 cubic metre

T = 700oC = 273 + 700 = 973 M

P = nRT / V =( n × 8.314 × 973)/(1.639 × 10-5)

Means pressure of CO2 depends on the number of moles.

In the case of helium also the volume and temperature are the same as above. Hence the volume of helium depend on its number of moles.

If the number of moles of CO2 and He are same then both will have the same pressure. Otherwise the pressure will be different.